Could you please help me understand how to work this problem: 1/3 log[7](8)-1/2 log[7](81)= log[7](x)
Answer by stanbon(75887)

1/3 log[7](8)-1/2 log[7](81)= log[7](x)
------
log7(8^(1/3)) - log7(81^(1/2)) = log7(x)
----
log7(2) - log7(9) = log7(x)
----
log7(2/9) = log7(x)
----
x = 2/9
================
Cheers,
Stan H.
==================